Potassium nitrate reacts with carbonic acid to produce potassium carbonate and nitric acid. A chemist adds 15.15 g of potassium nitrate to 0.228 mol of carbonic acid, and gets 8.45 g potassium carbonate. Write a balanced chemical equation, including all the states of matter.  Determine the limiting reactant and the amount of potassium carbonate (in grams) produced. What is the theoretical yield of potassium carbonate? Determine the percent yield of potassium carbonate in this reaction


1
Expert's answer
2021-06-15T05:13:29-0400

2KNO3 + H2CO3 = K2CO3 + 2HNO3

According to the equation, n (KNO3) = 2 x n (H2CO3) = 2 x n (K2CO3)

n = m / M

M (KNO3) = 101.1 g/mol

M (H2CO3) = 62.03 g/mol

M (K2CO3) = 138.2 g/mol

n (K2CO3) = 8.45 / 138.2 = 0.061 mol

n (KNO3)available = 15.15 / 101.1 = 0.15 mol

n (H2CO3) = 0.228 mol

The required amount of n (KNO3) = 2 x n (H2CO3) = 2 x 0.228 = 0.456 mol. This is more than the available amount (0.15 mol). That is why KNO3 is the limiting reactant. Due to this, the anticipated amount of K2CO3 is:

n (K2CO3) = 1/2 x n (KNO3) = 0.15 /2 = 0.075 mol

m (K2CO3) = n x M = 0.075 x 138.2 = 10.365 g

% (K2CO3) = 8.45 / 10.365 x 100 = 81.5%